# Design and Performance Simulation of the Electromagnetic Calorimeter at EicC

### v2

@@ 3.4 Design Optimization
The calorimeter design was optimized through detailed simulations, with particular emphasis on energy resolution, which is strongly influenced by the longitudinal radiation length and sampling structure. For the pCsI module, performance can be tuned by adjusting the total crystal length, while the Shashlik module allows further flexibility by varying both the number and thickness of its alternating absorber–scintillator layers. In addition, the lateral size of the module plays a crucial role in determining position resolution, whereas the impact of energy resolution on position resolution is relatively minor.
@@ 4 Neutral Pion Reconstruction
Another essential function of the ECAL is the reconstruction of mesons. Since decays into two photons in about 99% of cases, detecting both photons and separating those with small opening angles is crucial. The precision and efficiency of reconstruction depend on multiple factors, including energy and angular resolutions, detector acceptance, and shower separation capability. These factors reflect the combined performance of both the calorimeter and the reconstruction algorithm.
The event distribution from EicC collisions has shown in the right panel of Fig. 2. Low-momentum events are abundant in the barrel region, while high-momentum events are concentrated in the ion-Endcap. These two regimes place different demands on detector design and reconstruction strategies.
@@
where is the angle between each photon and the momentum direction in the laboratory frame after Lorentz transformation, and is the photon emission angle in the rest frame (taken as to yield the minimal opening angle). The Fig. 11 shows the distribution of the opening angle between the two decay photons for different momenta, obtained from Geant4 simulations, which is found to be consistent with Eq. (4.1).
In the relativistic limit, where the momentum is much greater than its rest mass (), Eq. (4.1) can be approximated by . This approximation clearly illustrates that the opening angle between the two photons decreases rapidly with increasing energy of , leading to a higher probability of cluster overlap in the calorimeter. This effect is particularly significant in the ion-Endcap region, where high-momentum mesons are predominant. To mitigate the overlap and improve the reconstruction efficiency, the calorimeter geometry can be optimized by increasing the distance between the ion-Endcap and the interaction point and by refining the cluster reconstruction algorithms.
Simulation results show there are three main factors can reduce the reconstruction efficiency. Firstly, the energy of each decay photon must exceed the electronic threshold to suppress noise and ensure reliable detection. Secondly, one of the photons may escape through the beam hole or hit the detector edge, resulting in a loss of geometric acceptance. This effect corresponds to the geometric acceptance efficiency, as illustrated in the left panel of Fig. 12. Thirdly, when the opening angle between the two photons is small, their electromagnetic showers may overlap, leading to the reconstruction of only a single cluster. This effect is associated with the cluster reconstruction efficiency, as shown in the right panel of Fig. 12, where high-momentum s exhibit lower reconstruction efficiency. Improving the angular resolution, for example by increasing the distance between the detector and the IP, together with enhancing reconstruction algorithm, can effectively mitigate this third source of inefficiency.
@@ 5 Conclusions
A comprehensive design and optimization study of the ECAL for the EicC has been presented, combining two complementary technologies to meet the distinct physics and geometric requirements of the detector: a pCsI crystal calorimeter in the e-Endcap and Shashlik style sampling calorimeters in the barrel and ion-Endcap regions.
Detailed Geant4 simulations, incorporating realistic detector responses, show that the proposed configuration meets the performance targets required by the EicC physics program. The pCsI modules offer excellent energy resolution of about , superior position resolution, and strong timing performance, benefiting from their high light yield. The Shashlik modules, optimized in sampling structure, radiation length, and light collection uniformity, provide energy and position resolutions of approximately and 5 mm at 1 GeV, respectively, along with effective electron–pion separation.
These results indicate that the current ECAL design can reliably support precision electron and photon measurements across the full acceptance of the EicC detector. Future work will focus on detailed digitization modeling, prototype construction and validation, and advancing the ECAL toward full engineering design and construction.

### v3

@@ 3.4 Design Optimization
The calorimeter design was optimized through detailed simulations, with particular emphasis on energy resolution, which is strongly influenced by the longitudinal radiation length and sampling structure. For the pCsI module, performance can be tuned by adjusting the total crystal length, while the Shashlik module allows further flexibility by varying both the number and thickness of its alternating absorber–scintillator layers. In addition, the lateral size of the module plays a crucial role in determining position resolution, whereas the impact of energy resolution on position resolution is relatively minor. In this section, the Edep-based method was employed for both the simulation and the reconstruction in order to study the intrinsic performance of different design configurations.
@@ 4 Neutral Pion Reconstruction
The reconstruction of neutral pions () is an important physics objective and a key performance benchmark of the ECAL. As a primary product of hadronization in deep inelastic scattering, measurements provide valuable information on parton fragmentation and nucleon structure, particularly in semi-inclusive and exclusive processes. In addition, the well-known invariant mass serves as a standard reference for ECAL calibration and energy resolution validation. Therefore, efficient and precise reconstruction is both a physics requirement and a crucial indicator of the overall ECAL performance.
Since decays into two photons in about 99% of cases, detecting both photons and separating those with small opening angles is crucial. The precision and efficiency of reconstruction depend on multiple factors, including energy and angular resolutions, detector acceptance, and shower separation capability. These factors reflect the combined performance of both the calorimeter and the reconstruction algorithm.
The distribution of mesons from EicC collisions was shown in the right panel of Fig. 2. Low-momentum events are abundant in the barrel region, while high-momentum events are concentrated in the ion-Endcap. These two regimes place different demands on detector design and reconstruction strategies.
@@
where is the angle between each photon and the momentum direction in the laboratory frame after Lorentz transformation, and is the photon emission angle in the rest frame. When , both photons are perpendicular to the flight direction, and in this case the opening angle reaches its minimum value:
Fig. 13 shows the distribution of the opening angle between the two decay photons for different momenta, obtained from Geant4 simulations, which is found to be consistent with Eq. (4.2). In the relativistic limit, where the momentum is much greater than its rest mass (), Eq. (4.2) can be approximated by . This approximation clearly illustrates that the opening angle between the two photons decreases rapidly with increasing energy of , leading to a higher probability of cluster overlap in the calorimeter. This effect is particularly significant in the ion-Endcap region, where high-momentum mesons are predominant. To mitigate the overlap and improve the reconstruction efficiency, the calorimeter geometry can be optimized by increasing the distance between the ion-Endcap and the interaction point and by refining the cluster reconstruction algorithms.
Simulation results using a more realistic geometry, as illustrated in the left panel of Fig. 4, indicate three main factors that can degrade the reconstruction efficiency. First, the energy of each decay photon must exceed the electronic threshold to suppress noise and ensure reliable detection. In this study, each photon cluster is required to satisfy a minimum energy threshold of MeV. Second, one of the photons may escape through the beam hole or hit the detector edge, resulting in a loss of geometric acceptance. The geometric acceptance efficiency is defined as the fraction of generated events for which both decay photons fall within the ECAL acceptance, as illustrated in the left panel of Fig. 14.
Third, when the opening angle between the two photons is small, their electromagnetic showers may overlap, leading to the reconstruction of a single cluster. The reconstruction efficiency is defined as the fraction of events in which two photon clusters are reconstructed in the ECAL and their invariant mass falls within the – MeV window, relative to the number of events satisfying the geometric acceptance requirement. As shown in the right panel of Fig. 14, high-momentum mesons exhibit lower reconstruction efficiency. Improving the angular resolution, for example by increasing the distance between the detector and the IP, together with optimizing the reconstruction algorithm, can effectively mitigate this inefficiency.
@@ 5 Conclusions
A comprehensive design and optimization study of the ECAL for the EicC has been presented, combining two complementary technologies to meet the distinct physics and geometric requirements of the detector: a pCsI crystal calorimeter in the e-Endcap and Shashlik-style sampling calorimeters in the barrel and ion-Endcap regions.
Detailed Geant4 simulations, incorporating realistic detector responses, show that the proposed configuration meets the performance targets required by the EicC physics program. The pCsI modules offer excellent energy resolution of , superior position resolution, and strong timing performance, benefiting from their enhanced effective light yield. The Shashlik modules, optimized in sampling structure, radiation length, and light collection uniformity, provide energy and position resolutions of and 5 mm at 1 GeV, respectively, along with effective electron–pion separation.
These results indicate that the current ECAL design can reliably support precision measurements of electrons and photons across the full acceptance of the EicC detector. Future work will focus on detailed digitization modeling, prototype construction and validation, and advancing the ECAL toward full engineering design and construction.